OpenStudy (anonymous):

the sum of 4/3 and four times is equal to 5/3 subtracted from five times the number HELP

OpenStudy (bahrom7893):

(4/3) + 4x = 5/3 - 5x

OpenStudy (anonymous):

and what do i do next? this is the first time i see this

OpenStudy (bahrom7893):

Sum of 4/3 and 4 times the number: let x be that number.. 4 times the number is 4x Sum of 4/3 and 4 times the number is: (4/3)+4x

OpenStudy (bahrom7893):

wait i made a mistake..

OpenStudy (bahrom7893):

(4/3) + 4x = 5x - (5/3)

OpenStudy (bahrom7893):

4/3 + 5/3 = 5x-4x 9/3 = x x = 3